# ———————————————————————————————————–

### v2

@@ Appendix B Proofs of Proposition 1 and Corollary 1
where ${\bf n}!:=n_{1}!\dots n_{p}!$ and $\mathcal{I}({\bf n};n)$ is the subset of locations ${\bf i}=\{i_{j,k}\}$ leading to non-overlapping patterns with counts set ${\bf n}$, such that all patterns are at distance of at least $M$ from each other, and $\omega({\bf i},{\bf n})\in\{0,1\}^{n}$ is the basis vector (trajectory) obtained by placing the all the patterns corresponding to the counts set ${\bf n}$ at locations prescribed by ${\bf i}$. More precisely ${\bf i}\in\mathcal{I}({\bf n};n)$ if for any two pairs $(j,k)$ and $(\tilde{j},\tilde{k})$ with $i_{j,k}\leq i_{\tilde{j},\tilde{k}}$ one has $i_{j,k}+|\alpha^{(j)}|+M-1<i_{\tilde{j},\tilde{k}}$. The basis vector $|\omega({\bf i},{\bf n})\rangle$ consist of zeros except patterns $\alpha^{(j)}$ written at positions $i_{j,k}$ for $j\in\{1,\dots,p\}$ and $k\in\{1,\dots,n_{j}\}$.
To show that the remainder term is $o(1)$ note that any term in the superposition $|{\bf n},n\rangle$ corresponding to a specific product of $\sigma$s is a vector of the standard basis, and any such vector $|\omega\rangle$ has at most a fixed number $M\cdot|{\bf n}|$ of $1$s. Note also that the action of applying a pattern $\sigma^{\alpha}_{i}$ to the vacuum cannot be reversed by applying subsequent $\sigma$s, since these contain only creation operators. This means that for a given $|\omega\rangle$, the locations ${\bf i}=\{i_{j,k}\}$ of the $\sigma$s producing this vector are limited to an area of size $2M$ around each $1$ in the sequence, or in other words, the coefficient of $|\omega\rangle$ is bounded by $(2M)^{M|{\bf n}|}$. On the other hand, the number of basis vectors $|\omega\rangle$ obtained by applying patterns such that at least two overlap or are at distance smaller than $M$ from each other is $o(n^{|{\bf n}|})$ since the number of possible locations for two such $\sigma$s is $O(n)$. Therefore, the remainder term in (29) is $o(1)$.
@@ Appendix C Proofs of Theorem 2 and Lemma 2
In order to prove the theorem, we will make use of the method of moments; this can be done since the moment problem corresponding to the moments of Poisson and Gaussian random variables admits a unique solution (this can be seen for instance using Cramér condition, see Lin (2017); Schmüdgen (2020)).
@@
Splitting of the sum based on non-overlapping groups of $\sigma$s. Let us consider a term $\sigma^{\alpha}_{i_{1}}\sigma^{\alpha*}_{j_{1}}\cdots\sigma^{\alpha}_{i_{r}}\sigma^{\alpha*}_{j_{r}}$ and represent each $\sigma^{\alpha}_{i}$ or $\sigma^{\alpha*}_{i}$ as a block of length $|\alpha|$ covering positions $\{i,i+1,\dots i+|\alpha|-1\}$ of the string $\{1,2,\dots n\}$. Depending on the overlapping pattern of the blocks, the indices can be split (uniquely) in a number $s=s(i_{1},j_{1},\dots,i_{r},j_{r})$ of groups ($1\leq s\leq 2r$) such that blocks in different groups do not overlap, and each group cannot be split into further non-overlapping sub-groups. Among these groups we identify $g=g(i_{1},j_{1},\dots,i_{r},j_{r})$ special groups characterised by the fact that they are made up of one or more pairs of blocks of the type $(\sigma_{z}^{\alpha*}\sigma^{\alpha}_{z})$ for some $z$. We call ${\cal P}_{0}$ the set of such groups. Note that not all groups associated to a product of $\sigma$s can be in $\mathcal{P}_{0}$ because the order of $\sigma_{z}^{\alpha*}$ and $\sigma_{z}^{\alpha}$ in the latter is opposite to that in which such terms appear in $N_{\alpha}(n)$.
For example, consider the pattern $\alpha=$$11$” for $n=10$ and the term
@@
Let us now look at the expected value of a given product of $\sigma$s. We have
@@
$\mathbf{1}^{\perp}:=\{x\in M(\mathbb{C}^{D}):\tilde{\rho}^{\rm ss}(x)=0\},$
@@
  • for every $m\geq 1$, the norm of ${\cal A}^{m}_{0}=\tilde{\mathcal{T}}^{m}$ restricted to $\mathbf{1}^{\perp}$ is less or equal than $c\lambda^{m}$ for some $c>0$ and $01>\lambda>$ 0.
@@
Recall that each term in (30) gives rise to a a certain set of groups of indices made up of overlapping blocks, and each group corresponds to a map ${\cal C}_{u,i}$ such that the expectation is expressed as in equation in (31). We will show that in limit of large $n$, the only terms which contribute to the $m$-th derivative are those coming from certain configurations with $s-g=m$ where $s$ is the number of groups and $g$ is the number of groups in $\mathcal{P}_{0}$.
Taylor approximation for a given set $({\cal C}_{u,1},\dots\cal C_{u,s})$ up to order $m=s-g$. Let us consider the sum of all the terms in Eq. (30) coming from all the correlations corresponding to a given sequence of maps $\boldsymbol{\mathcal{C}}:=({\cal C}_{u,1},\dots,{\cal C}_{u,s})$. This contribution is given by the product between a combinatorial factor (independent of $n$) counting how many products of $\sigma$s in (30) lead to the same set of maps $\boldsymbol{\mathcal{C}}$ and the following sum
$\begin{split}&\frac{1}{n^{r}}\sum_{x_{0}+\dots+x_{s}=n-K}\tilde{\rho}^{\rm ss}_{u}\left({\cal C}_{u,1}\tilde{\mathcal{T}}^{x_{1}}_{u}\cdots\tilde{\mathcal{T}}^{x_{s-1}}_{u}{\cal C}_{u,s}(\mathbf{1})\right),\\ \end{split}$ (34)
@@
Let us now consider the Taylor expansion of the correlations in (34). The 00-th order term is 00 because at least one ${\cal C}_{i}$ is not in ${\cal P}_{0}$ and hence $\tilde{\rho}^{\rm ss}$ annihilates the result. This is because on one hand, $C_{i*}(\tilde{\rho}^{\rm ss})=\tilde{\rho}^{\rm ss}$ for all $C_{i}\in\mathcal{P}_{0}$, but on the other hand at least one ${\cal C}_{i}$ is not in $\mathcal{P}_{0}$ and hence it contains a term $\mathcal{A}_{\pm 1}$ or $\mathcal{B}_{1}$ for which $\mathcal{A}_{\pm 1*}(\tilde{\rho}^{\rm ss})=\mathcal{B}_{1*}(\tilde{\rho}^{\rm ss})=0$ since $\tilde{K}_{1}{|{\chi^{\rm ss}}\rangle}=0$.
@@
The key additional ingredient is the fact that for $\mathcal{C}_{i}\notin\mathcal{P}_{0}$, the factors $\tilde{\mathcal{T}}^{x_{i}}\mathcal{C}_{i}$ are exponentially decreasing, cf. equation (32). This will provide more conservative upper bounds for the derivatives of the sum (34), as detailed below. Before considering that, note that the exponential bound can also be used to get an alternative proof of the convergence to 00 of
@@
    Hence differentiating $\tilde{\mathcal{T}}^{x_{i-1}}_{u}$ for which $\mathcal{C}_{u,i}$ is not in $\mathcal{P}_{0}$ removes the exponential decay associated to un-differentiated term $\tilde{\mathcal{T}}^{x_{i-1}}_{u}\mathcal{C}_{u,i}$ so that the upper bound for this contribution to the derivative is $n^{g+1}n^{-1/2}n^{-r}$. Since $g<r$ we have $g+1-1/2-r<0$ so the contribution to the derivative decays (even though it decays slower than the previous sub-sums, which will be important when considering higher derivatives).
@@
Recall that we are considering derivatives up to order $m=s-g$ which is equal to the number of $\mathcal{C}_{u,i}$ not in $\mathcal{P}_{0}$. According to the argument above, the most favourable positions for the derivatives is on different terms $\mathcal{T}^{x_{i-1}}_{u}\mathcal{C}_{u,i}$. Therefore, for a derivative of order $k\leq m$ the entire derivative can be upper bounded as
@@
Since $k\leq m=s-g$ the exponent is smaller than $t:=-r+g+(s-g)/2$. We will show that $t\leq 0$ with equality if and only if $m=s-g$ and the configuration $\boldsymbol{\mathcal{C}}$ is such that all groups in $\mathcal{P}_{0}$ consists of 2 perfectly overlapping $\sigma$s and the groups not in $\mathcal{P}_{0}$ consists of single blocks. Indeed for given total number $2r$ of $\sigma$ blocks, $s+g$ is maximum if all groups in $\mathcal{P}_{0}$ have just two $\sigma$ blocks and all others have a single block, so $2g+s-g=s+g\leq 2r$, which implies $t\leq 0$. In conclusion, for any configuration $\boldsymbol{\mathcal{C}}$, the derivatives below order $m=s-g$ decay asymptotically, and that of order $m$ does as well unless $\boldsymbol{\mathcal{C}}$ is of the special type described above. Derivatives of order above $m$ will be treated below. Note that for the special configurations, $m$ is even so all odd derivatives decay, as is expected from the fact that the intensity of the limit Poisson distribution is proportional to $u^{2}$. We will now compute the limit of the $m$th derivative for the special configuration.
If $\mathcal{C}_{i}\in\mathcal{P}_{0}$ consists of two blocks of $\sigma$s then the corresponding product is
@@
Finally, we need to compute the combinatorial factor counting the number of terms in the expectation which produce the desired set $\boldsymbol{\mathcal{C}}$ consisting of $g=r-m/2$ groups of two $\sigma$s in $\mathcal{P}_{0}$ plus $m$ groups of single $\sigma$s (with $s=r+m/2$ total number of groups).
We will show that this is exactly $S(r,m/2)$ (Stirling number of second type) i.e. the number of partitions of a set of $r$ elements into $m/2$ non-empty subsets. For this it is enough to show that the numbers of ways in which we can pair the $\sigma$s is in a bijection with the partitions in $m/2$ classes of $r$ elements. Consider the collection of $\sigma$s
@@
Reminder of the Taylor approximation (order $s-g+1$). The last thing we need to check is that the remainder corresponding to the sum of all the terms in Eq. (30) with $s$ non-overlapping blocks is negligible; the reminder is given by
@@
First suppose that $s-g>0$. We can prevent the maps which do not belong to ${\cal P}_{0}$ to be annihilated by $\rho$ and to cause an exponential decay differentiating and repeating the same computations as in the case of number operators. If we consider the $m$-th term in the Taylor expansion up to $m=s-g$, we can see that it grows at most as $n$ to the power $g+m/2-r/2$; if we want the exponent to be bigger or equal than 00, we need that $m\geq r-2g$; since $m\leq s-g$, one realises that $g+m/2-r/2$ can at most be equal to 00 and this is true when $m=s-g$ and $s+g=r$, which means that every group in ${\cal P}_{0}$ is of the form $\tilde{\sigma}_{z}^{\alpha*}(u)\tilde{\sigma}^{\alpha}_{z}(u)$ and all the other groups are singletons. However, if a group is composed by a single element, since we centered the random process, the first derivative will not be enough to cancel the exponential decay. Therefore, if $s-g>0$, the corresponding terms will decrease to 00.
On the other hand, if $r$ is even and $s=g=r/2$, the 00-th order term is equal to $1$, hence the leading term comes from the case when $r$ is even and $g=r/2$. In this case one can see that the limit quantity is equal to $1$ times the way we can pair the $\tilde{\sigma}(u)$’s in groups of the type $\tilde{\sigma}_{z}^{\alpha*}(u)\tilde{\sigma}^{\alpha}_{z}(u)$ and this is given by $(r-1)!!$! (which is the number of partition into pairs of a set of $r$ elements).
@@
~$\tilde{\mathcal{R}}_{*}\dot{\tilde{\mathcal{T}}}_{*}(\tilde{\rho}^{\rm ss})$ $=$ ${|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}$ (36)
@@
$Y$ ::=:= ${|{(\tilde{K}_{1}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}+\dot{\tilde{K}}_{1})\chi^{\rm ss}}\rangle}$ (37)
@@
$\begin{split}&\mbox{Tr}(({\rm Id}+\tilde{\mathcal{R}}{\cal B}_{1})({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|})Y)=\\ &\langle\chi^{\rm ss}|Y\chi^{\rm ss}\rangle+\langle\chi^{\rm ss}|{\cal B}_{1*}\tilde{\mathcal{R}}_{*}(Y-\mbox{Tr}(Y){|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|})\chi^{\rm ss}\rangle=\\ &\langle\chi^{\rm ss}|Y\chi^{\rm ss}\rangle+\langle\chi^{\rm ss}|\tilde{\mathcal{T}}_{*}\tilde{\mathcal{R}}_{*}(Y-\mbox{Tr}(Y){|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|})\chi^{\rm ss}\rangle-\\ &\langle\chi^{\rm ss}|\mathcal{R}_{*}(Y-\mbox{Tr}(Y){|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|})\chi^{\rm ss}\rangle=\\ &\langle\chi^{\rm ss}|Y\chi^{\rm ss}\rangle-\langle\chi^{\rm ss}|(Y-\mbox{Tr}(Y){|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|})\chi^{\rm ss}\rangle=\mbox{Tr}(Y).\\ \end{split}$
@@ Appendix D Proof of Corollary 2
$-\Re(\langle\chi^{\rm ss},2\dot{\tilde{K}}_{0}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}+\ddot{\tilde{K}}_{0}\chi^{\rm ss}\rangle)$
@@
$-\Re(\langle\chi^{\rm ss},\ddot{\tilde{K}}_{0}\chi^{\rm ss}\rangle)=\sum_{k=0}\mbox{Tr}(\rho^{\rm ss}\dot{K}^{*}_{k}\dot{K}_{k}).$
@@ Appendix E Proof of Theorem 3 and Proposition 2
Let us consider a fixed ordered sequence of excitation patterns $\alpha^{(1)},\dots,\alpha^{(k)}$ (here we do not require them to be distinct). For an observation time $n$ big enough, the probability of observing such a sequence of patterns separated one from the other by more than $n^{\gamma}$ consecutive 00s is given by
@@
  • for any $k\geq 1$, for some constants $C>0$, $10<\lambda<$ 1, one has $\|{\cal E}_{0}^{k}\|\leq C\lambda^{k}$,
@@
On the other hand, by differentiating a block of the form ${\cal B}^{x}_{u,0}{\cal B}_{u,\alpha}$ twice, one stops the exponential decay and obtains a linear growth in $n$; however, we need to remember that every time we use a derivative, everything gets multiplied by $u/\sqrt{n}$, which is of the order $n^{-1/2+\epsilon^{\prime}}$. Therefore, if we consider the terms in the Taylor expansion up to order $2k$, the only one that does not decay to 00 is the one with $2k$ derivatives where we use two of them on the block $\tilde{\rho}^{\rm ss}({\cal B}^{x_{1}}_{u,0}{\cal B}_{u,\alpha^{(1)}}(\cdot))$ at the beginning and other two on each following block of the form ${\cal B}_{u,0}^{x_{i}}{\cal B}_{u,\alpha^{(i)}}(\cdot)$. Any other term where we have less derivatives involved or where we spend them in a different way is either 00 or decays at least as (for $k\geq 2$)
@@
uniformly in $u$ for $|u|\leq n^{\epsilon^{\prime}}$; indeed, for any derivative used in a different way (they cannot be more than $2(k-1)$ because two of them need to be used for the first block), one gains a growth of $n^{1/2}$, but suffers a decay of at least $n^{\gamma}\lambda^{n^{\gamma}}$.
@@
$\begin{split}&\tilde{\rho}^{\rm ss}(\dot{\cal B}_{0}{\cal R}_{0}\dot{\cal B}_{\alpha}({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}))=\\ &2\Re\left(\langle\tilde{K}_{\alpha_{|\alpha|}}\cdots\tilde{K}_{\alpha_{2}}\dot{\tilde{K}}_{1}\chi^{\rm ss}|\chi^{\rm ss}\rangle\times\right.\\ &\left.\langle\chi^{\rm ss}|\tilde{K}_{\alpha_{|\alpha|}}\cdots\tilde{K}_{\alpha_{1}}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}\rangle\right).\\ \end{split}$
@@
where by ${\cal R}_{0*}$ we mean the Moore-Penrose inverse and ”${\rm rem}$” is a term which is gives 00 when evaluated against ${\cal B}_{\alpha}({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}).$ Equivalently,
(¨$\begin{split}&(\ddot{{\cal B}}_{0*}+2\dot{\cal B}_{0*}{\cal R}_{0*}\dot{\cal B}_{0*})(\tilde{\rho}^{\rm ss})=\\ &2({\rm Id}-{\cal B}_{0*})({|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|})+{\rm rem}^{\prime},\end{split}$
@@
$\begin{split}&({\rm Id}-{\cal B}_{0*})({|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|})=\\ &{|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|}-\\ &{|{\tilde{K}_{0}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{\tilde{K}_{0}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|}.\\ \end{split}$
@@
$\begin{split}&({\rm Id}-{\cal B}_{0*})({|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|})=\\ &{|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{\dot{\tilde{K}}_{0}\chi^{\rm ss}}|}+{|{\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|}-\\ &{|{\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{\dot{\tilde{K}}_{0}\chi^{\rm ss}}|}.\\ \end{split}$
@@
~$\tilde{{\cal B}}_{\alpha,u}=e^{\lambda_{tot}|\alpha|u^{2}/n}{\cal B}_{\alpha,u}$
@@
Let us first point out some properties of the maps ~$\tilde{{\cal B}}_{\alpha,u}$:
@@
    ~$\tilde{{\cal B}}_{0,u}(\cdot)=a(u)l(u)(\cdot)r(u)+\tilde{{\cal E}}_{0,u}(\cdot)$
@@
    • ~$\tilde{{\cal E}}_{0,u}(r(u))=0$ and
    • ~$\|\tilde{{\cal E}}^{k}_{0,u}\|\leq C\lambda^{k}$ for some $C\geq 0$ and $10<\lambda<$ 1.
    The order of the reminder in the expression $a(u)$ is due to the fact that $\dot{\lambda}$ is equal to the first derivative at 00 of the spectral radius of ${\cal B}_{0,u}$ which is equal to 00 because it attains a maximum there; the fact that $\ddot{\lambda}=0$ as well is due to the multiplicative factor in front of ${\cal B}_{0,u}$ in the definition of ~$\tilde{{\cal B}}_{0,u}$.
@@
    Indeed, the first derivative of $l(u)(\tilde{{\cal B}}_{\alpha,u}(\cdot))$ at 00 is given by

### v3

@@ Appendix B Proofs of Proposition 1 and Corollary 1
[…]
@@ Appendix C Proofs of Theorem 2 and Lemma 2
In order to prove the theorem, we will make use of the method of moments; this can be done since the moment problem corresponding to the moments of Poisson and Gaussian random variables admits a unique solution (this can be seen for instance using Cramér condition, see [140, 141]).
@@
[…]
For example, consider the pattern $\alpha=$1111” for $n=10$ and the term
@@
[…]
@@
$\mathbf{1}^{\perp}:=\{x\in B(\mathcal{H}_{sa}):\tilde{\rho}^{\rm ss}(x)=0\},$
@@
  • for every $m\geq 1$, the norm of ${\cal A}^{m}_{0}=\tilde{\mathcal{T}}^{m}$ restricted to $\mathbf{1}^{\perp}$ is less or equal than $c\lambda^{m}$ for some $c>0$ and $1>\lambda>0$.
@@
Recall that each term in (32) gives rise to a a certain set of groups of indices made up of overlapping blocks, and each group corresponds to a map ${\cal C}_{u,i}$ such that the expectation is expressed as in equation in (33). We will show that in limit of large $n$, the only terms which contribute to the $m$-th derivative are those coming from certain configurations with sg=ms-g=m where $s$ is the number of groups and $g$ is the number of groups in $\mathcal{P}_{0}$.
[…]
@@
Let us now consider the Taylor expansion of the correlations in (36). The 0-th order term is 0 because at least one ${\cal C}_{i}$ is not in ${\cal P}_{0}$ and hence $\tilde{\rho}^{\rm ss}$ annihilates the result. This is because on one hand, $C_{i*}(\tilde{\rho}^{\rm ss})=\tilde{\rho}^{\rm ss}$ for all $C_{i}\in\mathcal{P}_{0}$, but on the other hand at least one ${\cal C}_{i}$ is not in $\mathcal{P}_{0}$ and hence it contains a term $\mathcal{A}_{\pm 1}$ or $\mathcal{B}_{1}$ for which $\mathcal{A}_{\pm 1*}(\tilde{\rho}^{\rm ss})=\mathcal{B}_{1*}(\tilde{\rho}^{\rm ss})=0$ since $\tilde{K}_{1}{|{\chi^{\rm ss}}\rangle}=0$.
@@
The key additional ingredient is the fact that for $\mathcal{C}_{i}\notin\mathcal{P}_{0}$, the factors $\tilde{\mathcal{T}}^{x_{i}}\mathcal{C}_{i}$ are exponentially decreasing, cf. equation (34). This will provide more conservative upper bounds for the derivatives of the sum (36), as detailed below. Before considering that, note that the exponential bound can also be used to get an alternative proof of the convergence to 0 of
@@
    Hence differentiating $\tilde{\mathcal{T}}^{x_{i-1}}_{u}$ for which $\mathcal{C}_{u,i}$ is not in $\mathcal{P}_{0}$ removes the exponential decay associated to un-differentiated term $\tilde{\mathcal{T}}^{x_{i-1}}_{u}\mathcal{C}_{u,i}$ so that the upper bound for this contribution to the derivative is $n^{g+1}n^{-1/2}n^{-r}$. Since $g<r$ we have g+11/2r<0g+1-1/2-r<0 so the contribution to the derivative decays (even though it decays slower than the previous sub-sums, which will be important when considering higher derivatives).
@@
Recall that we are considering derivatives up to order m=sgm=s-g which is equal to the number of $\mathcal{C}_{u,i}$ not in $\mathcal{P}_{0}$. According to the argument above, the most favourable positions for the derivatives is on different terms $\mathcal{T}^{x_{i-1}}_{u}\mathcal{C}_{u,i}$. Therefore, for a derivative of order $k\leq m$ the entire derivative can be upper bounded as
@@
[…]
@@
[…]
@@
Reminder of the Taylor approximation (order sg+1s-g+1). The last thing we need to check is that the remainder corresponding to the sum of all the terms in Eq. (32) with $s$ non-overlapping blocks is negligible; the reminder is given by
@@
First suppose that sg>0s-g>0. We can prevent the maps which do not belong to ${\cal P}_{0}$ to be annihilated by $\rho$ and to cause an exponential decay differentiating and repeating the same computations as in the case of number operators. If we consider the $m$-th term in the Taylor expansion up to m=sgm=s-g, we can see that it grows at most as $n$ to the power g+m/2r/2g+m/2-r/2; if we want the exponent to be bigger or equal than 0, we need that $m\geq r-2g$; since $m\leq s-g$, one realises that g+m/2r/2g+m/2-r/2 can at most be equal to 0 and this is true when m=sgm=s-g and $s+g=r$, which means that every group in ${\cal P}_{0}$ is of the form $\tilde{\sigma}_{z}^{\alpha*}(u)\tilde{\sigma}^{\alpha}_{z}(u)$ and all the other groups are singletons. However, if a group is composed by a single element, since we centered the random process, the first derivative will not be enough to cancel the exponential decay. Therefore, if sg>0s-g>0, the corresponding terms will decrease to 0.
On the other hand, if $r$ is even and $s=g=r/2$, the 0-th order term is equal to $1$, hence the leading term comes from the case when $r$ is even and $g=r/2$. In this case one can see that the limit quantity is equal to $1$ times the way we can pair the $\tilde{\sigma}(u)$’s in groups of the type $\tilde{\sigma}_{z}^{\alpha*}(u)\tilde{\sigma}^{\alpha}_{z}(u)$ and this is given by (r1)!!(r-1)!! (which is the number of partition into pairs of a set of $r$ elements).
@@
$\tilde{\mathcal{R}}_{*}\dot{\tilde{\mathcal{T}}}_{*}(\tilde{\rho}^{\rm ss})$ $=$ ${|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}$ (38)
@@
$Y$ $:=$ ${|{(\tilde{K}_{1}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}+\dot{\tilde{K}}_{1})\chi^{\rm ss}}\rangle}$ (39)
@@
[…]
@@ Appendix D Proof of Corollary 2
$\Re(\langle\chi^{\rm ss},2\dot{\tilde{K}}_{0}(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}+\ddot{\tilde{K}}_{0}\chi^{\rm ss}\rangle)$
@@
$\Re(\langle\chi^{\rm ss},\ddot{\tilde{K}}_{0}\chi^{\rm ss}\rangle)=\sum_{k=0}\mbox{Tr}(\rho^{\rm ss}\dot{K}^{*}_{k}\dot{K}_{k}).$
@@ Appendix E Proof of Theorem 3 and Proposition 2
Let us consider a fixed ordered sequence of excitation patterns $\alpha^{(1)},\dots,\alpha^{(k)}$ (here we do not require them to be distinct). For an observation time $n$ big enough, the probability of observing such a sequence of patterns separated one from the other by more than $n^{\gamma}$ consecutive 0s is given by
@@
  • for any $k\geq 1$, for some constants $C>0$, $0<\lambda<1$, one has $\|{\cal E}_{0}^{k}\|\leq C\lambda^{k}$,
@@
On the other hand, by differentiating a block of the form ${\cal B}^{x}_{u,0}{\cal B}_{u,\alpha}$ twice, one stops the exponential decay and obtains a linear growth in $n$; however, we need to remember that every time we use a derivative, everything gets multiplied by $u/\sqrt{n}$, which is of the order $n^{-1/2+\epsilon^{\prime}}$. Therefore, if we consider the terms in the Taylor expansion up to order $2k$, the only one that does not decay to 0 is the one with $2k$ derivatives where we use two of them on the block $\tilde{\rho}^{\rm ss}({\cal B}^{x_{1}}_{u,0}{\cal B}_{u,\alpha^{(1)}}(\cdot))$ at the beginning and other two on each following block of the form ${\cal B}_{u,0}^{x_{i}}{\cal B}_{u,\alpha^{(i)}}(\cdot)$. Any other term where we have less derivatives involved or where we spend them in a different way is either 0 or decays at least as (for $k\geq 2$)
@@
uniformly in $u$ for $|u|\leq n^{\epsilon^{\prime}}$; indeed, for any derivative used in a different way (they cannot be more than 2(k1)2(k-1) because two of them need to be used for the first block), one gains a growth of $n^{1/2}$, but suffers a decay of at least $n^{\gamma}\lambda^{n^{\gamma}}$.
@@
[…]
@@
where by ${\cal R}_{0*}$ we mean the Moore-Penrose inverse and $\rm rem$}” is a term which is gives 0 when evaluated against ${\cal B}_{\alpha}({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}).$ Equivalently,
$\begin{split}&(\ddot{{\cal B}}_{0*}+2\dot{\cal B}_{0*}{\cal R}_{0*}\dot{\cal B}_{0*})(\tilde{\rho}^{\rm ss})=\\ &2({\rm Id}-{\cal B}_{0*})({|{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}\rangle}{\langle{(\mathbf{1}-\tilde{K}_{0})^{-1}\dot{\tilde{K}}_{0}\chi^{\rm ss}}|})+{\rm rem}^{\prime},\end{split}$
@@
[…]
@@
[…]
@@
$\tilde{{\cal B}}_{\alpha,u}=e^{\lambda_{tot}|\alpha|u^{2}/n}{\cal B}_{\alpha,u}$
@@
Let us first point out some properties of the maps $\tilde{{\cal B}}_{\alpha,u}$:
@@
    $\tilde{{\cal B}}_{0,u}(\cdot)=a(u)l(u)(\cdot)r(u)+\tilde{{\cal E}}_{0,u}(\cdot)$
@@
    • $\tilde{{\cal E}}_{0,u}(r(u))=0$ and
    • $\|\tilde{{\cal E}}^{k}_{0,u}\|\leq C\lambda^{k}$ for some $C\geq 0$ and $0<\lambda<1$.
    The order of the reminder in the expression $a(u)$ is due to the fact that $\dot{\lambda}$ is equal to the first derivative at 0 of the spectral radius of ${\cal B}_{0,u}$ which is equal to 0 because it attains a maximum there; the fact that $\ddot{\lambda}=0$ as well is due to the multiplicative factor in front of ${\cal B}_{0,u}$ in the definition of $\tilde{{\cal B}}_{0,u}$.
@@
    Indeed, the first derivative of $l(u)(\tilde{{\cal B}}_{\alpha,u}(\cdot))$ at 0 is given by